### v1

@@ 2.2 Non-dimensional numbers
The most basic non-dimensional number related to the transport of $\mathrm{H_{2}}$ is the Sherwood number. The production rate of hydrogen flux is constant at the electrode, and $J$ is a constant value in time. The transport of hydrogen results in a surface-averaged concentration, $\bar{c_{e}}$ at the electrode surface. To illustrate the transport, we compare the $\bar{c_{e}}$ with the concentration value in the bulk liquid, for which we adopt the top of the domain. Since the electrolyte solution is initially saturated ($\zeta=1$), the initial value $c_{0}=c_{sat}$, it yields the Sherwood number of production of hydrogen,
$\overline{\mathrm{Sh_{e}}}=\frac{JL_{ref}}{D(\bar{c_{e}}-c_{sat})}.$ (5)
The bar symbol is used to mark the surface-averaged response parameters. By introducing the boundary layer thickness $\delta_{\mathrm{H_{2}}}=D\Delta c/J$, this Sherwood number could be expressed as $\mathrm{Sh}=D_{b}/\delta_{\mathrm{H_{2}}}$. For pure diffusion, the $\delta_{\mathrm{H_{2}}}$ ultimately reaches the electrolysis cell height irrespective of the current density so that the Sherwood could maintain stability for all the cases without the effect of the bubbles.
For the mass transfer considering local volume change, the hydrogen transport to the bubble will be considered. This leads to the expression for the bubble Sherwood number,
[…]
where $A_{\Sigma}$ is the interface surface area, $4\pi R$, and the concentration difference between the electrode and bubble interface, $c_{e}-c_{\Sigma}$. According to eq. 31, the concentration of the interface $c_{\Sigma}=c_{sat}$. Since we assume the configuration with an ambient temperature and pressure, the mass transfer rate across the interface can be defined as:
[…]
where $\mathcal{R},T_{0},P_{0}$ are the universal gas constant, ambient temperature, and pressure, respectively. Then it leads to a rather simple expression of bubble Sherwood number,
$\overline{\mathrm{Sh_{e}}}=\frac{P_{0}}{\mathcal{R}T_{0}}\frac{2R}{D(\bar{c_{e}}-c_{\Sigma})}\frac{dR}{dt}$ (8)
Apart from this crucial Sherwood number, which measures the mass transfer, other nondimensional numbers are also vital as control parameters for physics. First, the Schmidt number compares momentum diffusion to mass diffusion,
@@
Where the $\nu_{c}$ is the kinematic viscosity of the electrolyte. For these configurations with gravity added, the dynamics of bubbles in a gravitational field can be described by the Galilei and Bond numbers. The Galilei number compares the gravitational and viscous forces and is defined as:
@@
The ratio of gravitational and capillary forces is estimated by the Bond number,
@@
2.3 Governing equation
A model of the microscopic process involves the solution of the two-phase incompressible Navier-Stokes equations with phase change, surface tension, gravity, and contact line dynamics on the wall. Due to the constant temperature assumption, no equation for the thermal energy is needed.
The gas phase is called $\Omega_{d}(t)$, for the disperse phase, and the liquid phase is called $\Omega_{c}(t)$, for the continous phase. These two subdomains are separated by an infinitely thin interface $\Sigma(t)$. The entire domain is given by $\Omega=\Omega_{d}(t)\cup\Omega_{c}(t)\cup\Sigma(t)$. The normal vector $n_{\Sigma}$ at the interface points into $\Omega_{d}(t)$.
@@ 2.3 Governing equation
$\begin{split}\partial_{t}(\rho\mathbf{u})+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})&=-\nabla p+\nabla\cdot(2\mu\mathbf{D})\\ &\quad+\rho\mathbf{g}\quad\text{in }\Omega\backslash\Sigma.\end{split}$ (13)
where the density $\rho$ and viscosity $\mu$ remain constant in $\Omega_{d}$ and $\Omega_{c}$. The eq. 12 is the continuity equation, where $\mathbf{u}$ represents the velocity field. In the balance of momentum eq. 13, $p$ is the static pressure, $\mathbf{D}$ is the deformation tensor, and $g$ represents body force, which is gravity acceleration in this system. So, $a$ is replaced by $g$ for the following sections. The eq. 12 and eq. 13 are valid everywhere in the domain except at the interface, where additional conditions are needed (Tryggvason et al., 2011). The continuity equation requires that the amount of mass that leaves one phase $\Omega_{d}(\Omega_{c})$ must be transferred to another phase $\Omega_{c}(\Omega_{d})$ since the infinitely thin interface region can not store any mass. It results in a jump condition across the interface,
$\rho(\mathbf{u}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}=0,$ (14)
where the jump notation has been introduced; $\mathbf{u_{\Sigma}}$ is the interface velocity and $\dot{m}$ is the mass transfer rate $\mathrm{kg/m^{2}s}$. The second jump condition is derived by applying the conservation of momentum to a control volume with an infinitely small thickness around the interface, and it reads,
$\mathbf{u}\otimes(\mathbf{u}-\mathbf{u_{\Sigma}})+p\mathbf{I}-2\mu\mathbf{D}\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}+\nabla_{\Sigma}\sigma.$ (15)
where $\mathbf{I}$ is the unit tensor, $\sigma$ is the surface tension and $k$ is the curvature of the interface. In the problem of hydrogen bubble growth, the interface is considered to have uniform surface tension, besides a no-slip boundary condition is applied at the interface, eq. 15 can be further simplified, see Fleckenstein and Bothe (2015).
$\left(p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}\right)=\sigma k\mathbf{n_{\Sigma}}$ (16)
The numerical method used in this work for interface transport is the Volume of Fluid (VOF) method, and is combined with one-fluid formulation of the governing equations. In the one-fluid approach, the jump condition eq. 14 and eq. 16 are replaced by source terms that act at the interface as singularities ($\delta$ function), a single set of Navier-Stokes equations for the entire domain $\Omega$ is solved.
@@
where the $\delta_{\Sigma}$ is the surface Dirac function which has a nonzero value only at the interface. So the system of eq. 17 eq. 18 is valid in the whole domain. Then, to determine the location of the interface, a marker function is required. The Heaviside function serves this purpose,
@@
The transport equation for Heaviside function $H(x,t)$ can be obtained from the following integral balance for a control volume $V$.
[…]
where the second term on the LHS represents the convective transport, and the last term is a source term that accounts for the mass transfer across the interface, which is null when $\dot{m}=0$. Converting the surface integral to volume integral, we can write it into differential form,
@@
eq. 21 is then served to compute the volume fraction, see section 2.5.
@@ 2.4 Concentration transport equation
This model can be used to compute the concentration field of the soluble hydrogen in a two-phase system with mass transfer by applying the two-scalar method of Fleckenstein and Bothe (2015). In the present study, we focus on pure incompressible gas bubbles, and we assume that no electrolyte species exists in the disperse phase (i.e., the electrolyte is not volatile). The system contains two different species: hydrogen (denoted with subscript 1) and electrolyte liquid (denoted with subscript 2). The overall mass transfer is entirely given by the transfer of the hydrogen species. It is worth pointing out that we do not need to solve the mass balance in the dispersed phase since no mixture exists inside the bubbles (pure hydrogen gas). The mass balance of hydrogen in the continous side $\Omega_{c}$ reads
$\partial_{t}\rho^{1}+\nabla\cdot(\rho^{1}\mathbf{u^{1}})=R,$ (22)
where $\rho^{1}$ is the partial density of hydrogen, $\mathbf{u^{1}}$ denotes the hydrogen velocity. $R$ is the reaction term that can model chemical reactions. It is coupled with the jump condition for the conservation of mass,
$\|\rho^{1}(\mathbf{u^{1}}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}||=||\dot{m}^{1}||=0.$ (23)
The average phase density and velocity of are derived from the respective species terms,
@@
where the superscript $2$ denotes the species electrolyte. Then the transport equation of hydrogen for incompressible flow can be written as
$\partial_{t}\rho^{1}+\mathbf{u}\cdot\nabla\rho^{1}+\nabla\cdot J^{1}=R,$ (26)
@@
The mass transfer rate of hydrogen can be derived from eq. 23,eq. 14.
$\begin{split}\dot{m}^{1}&=\rho^{1}(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}+\rho^{1}(\mathbf{u}^{1}-\mathbf{u})\cdot\mathbf{n}_{\Sigma}\\ &=\frac{\rho^{1}}{\rho_{c}}\cdot\dot{m}+J^{1}\cdot\mathbf{n}_{\Sigma}\end{split}$ (28)
@@
where the $D^{1}$ is the hydrogen diffusion coefficient. Combining the eq. 29 and eq. 28, the mass transfer rate of hydrogen reads
$\dot{m}=-\frac{M^{1}D^{1}}{1-\frac{\rho^{1}}{\rho_{c}}}\frac{\partial c^{1}}{\partial\mathbf{n_{\Sigma}}},$ (30)
where the molar concentration has been introduced, i.e. $c^{1}=\rho^{1}/M^{1}$ and $M^{1}$ is the molar mass. It gives the mass transfer rate evaluated from the continuous side of the interface, where the concentration of the dissolved species is generally variable. Since we have only hydrogen, which can be transferred between phases, in the following, we will omit the species indicator as we will always refer to the concentration of the soluble hydrogen in the liquid phase $\Omega_{c}$.
To compute the concentration gradient, we need the hydrogen concentration at the liquid side of the interface. For a gas-liquid system at equilibrium, we can employ Henry’s law to compute the concentration on the liquid side of the interface.
$(c_{c})_{\Sigma}=\frac{(c_{d})_{\Sigma}}{\mathrm{He}},$ (31)
where $\mathrm{He}$ is the Henry coefficient, and it is taken as a constant for the present work. So, the hydrogen concentration at the liquid side of the interface $(c_{c})_{\Sigma}$ is immediately computed. $(c_{d})_{\Sigma}=\rho_{d}/M$ is a constant as the density should be equal everywhere inside the bubble ($\Omega_{d}$).
For numerical integration, we rewrite the mass balance eq. 26 for the hydrogen diffusion process. As we discussed, the mass balance will be done in the continuous region $\Omega_{c}$ and for a non-reactive flow ($R=0$). Since the hydrogen dissolved in $\Omega_{c}$ is the only species that we need for mass transportation. The phase indicator will be omitted in the following, i.e., $c=c_{c}$.
[…]
Combine the jump condition eq. 23, the final differential form could be deduced.
@@ 2.5 Numerical methodology
The phase change model is derived from the work of Gennari et al. (2022). The governing equation shown in section 2.3 is solved by open source solver Basilisk (HTTP://basilisk.fr/) which is a finite-volume P.D.E. solver on adaptive Cartesian grids. Using adaptive mesh refinement (AMR) in regions with large gradients makes the approach particulally suitable for multiscale processes such as interfacial flows. In interfacial flows, a fine mesh is typically required around the gas-liquid interface but not the entire domain. The shape of the domain is always a square $L_{0}\times L_{0}$ in axisymmetric configuration (cube $L_{0}\times L_{0}\times L_{0}$ in 3D). The grid is organized following a hierarchical quadtree (octree) structure, where each cell can be further divided into four child cells (eight in 3D), and a level is assigned to each cell according to the relative tree structure. The root cell is at level 0, and its size $\Delta$ is the same as the whole numerical domain ($\Delta=L_{0}$). A generic cell at level $l$ has a resolution of $\Delta(l)=\frac{L_{0}}{2^{l}}$. The grid structure in Basilisk allows neighboring cells to vary by up to one level, meaning each cell edge/face can communicate with no more than two finer edges/faces. An example of a quadtree grid with different levels of refinement and the relative tree structure is shown in fig. 4.
Figure 4: An example of quadtree discretisation (left) together with its logical representation (right). The level of the cells in the tree is also given (Popinet, 2012).
The VOF method is one of the most widely used numerical approaches for the modeling of two-phase immiscible fluids. The starting point for the derivation of the VOF approach is the one-fluid formulation presented in section 2.3, which is vital to the phase change problem. The volume fraction of the continuous phase is defined as
@@
$f_{c}=\begin{cases}0&\text{if the cell is pure gas},\\ 1&\text{if the cell is pure liquid},\\ [0,1]&\text{if the cell is mixed}.\end{cases}$ (35)
The volume fraction of the dispersed phase $f_{d}$ is implicitly described by the relationship $f_{c}+f_{d}=1$. So, only one transport equation of H needs to be solved. By applying the incompressibility constraint, the integrated form of eq. 21 is
[…]
According to the finding of Scardovelli and Zaleski (1999), the integration of the transport eq. 36 is performed in two steps, namely the reconstruction step and the propagation step. First, the interface is approximated with a line/plane in each interfacial cell. Second, the fluxes of volume fraction across the cell boundaries are computed, and eq. 36 is integrated in time. The geometric reconstruction of the interface is based on the piecewise linear interface construction (PLIC) method, where the interface is approximated as a line (plane) in axisymmetric configuration (3D).

### v4

@@ 2.2 Non-dimensional numbers
The most basic non-dimensional number related to the transport of $\mathrm{H_{2}}$ is the Sherwood number. The production rate of hydrogen flux is constant at the electrode, and $J$ is a constant value in time. The transport of hydrogen results in a surface-averaged concentration, $\bar{c_{e}}$ at the electrode surface. To illustrate the transport, we compare $\bar{c_{e}}$ with the bulk liquid concentration, defined here as the concentration at the top of the domain. Given that the electrolyte solution is initially saturated ($\zeta=1$), and $c_{0}=c_{s}$, this yields the following Sherwood number for hydrogen production,
$\overline{\mathrm{Sh_{e}}}=\frac{JD_{b}}{D(\bar{c_{e}}-c_{s})}.$ (5)
The bar symbol is used to mark the surface-averaged response parameters. Introducing the boundary layer thickness $\delta_{\mathrm{H_{2}}}=D(\bar{c_{e}}-c_{s})/J$, this Sherwood number can be expressed as $\overline{\mathrm{Sh_{e}}}=D_{b}/\delta_{\mathrm{H_{2}}}$. The boundary layer thickness is, in principle, time dependent, so the Sherwood number also is time dependent. However, after several bubble detachments or in the presence of externally forced flow, the boundary layer thickness would eventually reach a statistical steady state, and a statistically steady Sherwood number will be reached.
For mass transfer accounting for local volume change, the Sherwood number of the bubble is expressed as follows to quantify the hydrogen transport across the interface,
$\overline{\mathrm{Sh_{b}}}=\frac{D_{b}\int_{\Sigma}\dot{m}\,ds}{A_{\Sigma}MD(\bar{c_{e}}-c_{s})},$ (6)
where $A_{\Sigma}=4\pi R^{2}$ is the interface surface area for a spherical bubble. And $\bar{c_{e}}-c_{s}$ represents the concentration difference between the electrode and bubble interface as the concentration of the interface $c_{\Sigma}=c_{s}$. Since we assume the configuration with an ambient temperature and pressure, the mass transfer rate across the interface can be defined as
$\int_{\Sigma}\dot{m}\,ds=\left(\frac{P_{0}}{\mathcal{R}T_{0}}\right)4\pi R^{2}\frac{dR}{dt}M,$ (7)
where $\mathcal{R},T_{0},P_{0}$ are the universal gas constant, ambient temperature, and pressure, respectively. Then it leads to a rather simple expression of the bubble Sherwood number,
$\overline{\mathrm{Sh_{b}}}=\frac{P_{0}}{\mathcal{R}T_{0}}\frac{2R}{D(\bar{c_{e}}-c_{\Sigma})}\frac{dR}{dt}.$ (8)
Apart from the Sherwood number, which measures the mass transfer, other important nondimensional numbers are the Schmidt, Galileo, and Bond numbers. First, the Schmidt number compares momentum diffusion to mass diffusion and is defined as
@@
where the $\nu_{c}$ is the kinematic viscosity of the electrolyte. The ratios of dominant forces acting on the bubble can be described using two additional dimensionless numbers: the Galileo and Bond numbers. The Galileo number, which compares gravitational to viscous forces, is given by
@@
When the bubble approaches detachment, gravitational effects become dominant and can be quantified by the Bond number, which represents the ratio of gravitational to capillary forces
@@
2.3 Fluid dynamical equations
A model of the microscopic process involves solving the two-phase incompressible Navier-Stokes equations with phase change, surface tension, gravity, and contact line dynamics on the wall. Due to the constant temperature assumption, no equation for the thermal energy is needed.
The gas phase is called $\Omega_{d}(t)$, for the disperse phase, and the liquid phase is called $\Omega_{c}(t)$, for the continuous phase. These two subdomains are separated by an infinitely thin interface $\Sigma(t)$. The entire domain is given by $\Omega=\Omega_{d}(t)\cup\Omega_{c}(t)\cup\Sigma(t)$. The normal vector $n_{\Sigma}$ at the interface points into $\Omega_{d}(t)$.
@@ 2.3 Fluid dynamical equations
$\begin{split}\partial_{t}(\rho\mathbf{u})+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})&=-\nabla p+\nabla\cdot(2\mu\mathbf{D})\\ &\quad+\rho\mathbf{g}\quad\text{in }\Omega\backslash\Sigma,\end{split}$ (13)
where the density $\rho$ and viscosity $\mu$ remain constant in $\Omega_{d}$ and $\Omega_{c}$. The Eq.˜12 is the continuity equation, where $\mathbf{u}$ represents the velocity field. In the balance of momentum Eq.˜13, $p$ is the static pressure, $\mathbf{D}$ is the deformation tensor, and $g$ represents body force, which is the gravitational acceleration in this system. So, $a$ is replaced by $g$ in the following sections. The Eq.˜12 and Eq.˜13 are valid everywhere in the domain except at the interface, where additional conditions are needed (Tryggvason_Scardovelli_Zaleski_2011). The continuity equation requires that the amount of mass that leaves one phase $\Omega_{d}(\Omega_{c})$ must be transferred to another phase $\Omega_{c}(\Omega_{d})$ since the infinitely thin interface region can not store any mass. It results in a jump condition across the interface,
$\|\rho(\mathbf{u}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}\|=\|\dot{m}\|=0,$ (14)
where the jump notation has been introduced(e.g. $\|\rho\|=\rho_{c}-\rho_{d}$); $\mathbf{u_{\Sigma}}$ is the interface velocity and $\dot{m}$ is the mass transfer rate $\mathrm{kg/m^{2}s}$. The second jump condition is derived by applying the conservation of momentum to a control volume with an infinitely small thickness around the interface, and it reads
$\|\rho\mathbf{u}\otimes(\mathbf{u}-\mathbf{u_{\Sigma}})+p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}+\nabla_{\Sigma}\sigma,$ (15)
where $\mathbf{I}$ is the unit tensor, $\sigma$ is the surface tension and $k$ is the curvature of the interface. In the problem of hydrogen bubble growth, the interface is considered to have uniform surface tension; besides, a no-slip boundary condition is applied at the interface. Eq.˜15 can be further simplified fleckenstein2015 in the form
$\|p\mathbf{I}-2\mu\mathbf{D}\|\cdot\mathbf{n_{\Sigma}}=\sigma k\mathbf{n_{\Sigma}}.$ (16)
The numerical method used in this work for interface transport is the Volume of Fluid (VOF) method, and is combined with a one-fluid formulation of the governing equations. In the one-fluid approach, the jump conditions Eq.˜14 and Eq.˜16 are replaced by source terms that act at the interface as singularities ($\delta$ function), a single set of Navier-Stokes equations for the entire domain $\Omega$ is solved as follows
@@
where the $\delta_{\Sigma}$ is the surface Dirac distribution. Therefore, the system of Eqs˜17 and 18 is valid for the whole domain. Then, to determine the location of the interface, a marker function is required. The Heaviside function serves this purpose
@@
The transport equation for Heaviside function $H(x,t)$ can be obtained from the following integral balance for a control volume $V$
$\begin{split}\int_{V}\partial_{t}H\,dV+\oint_{\partial V}H\mathbf{u}\cdot\mathbf{n}\,dS\\ +\int_{\Sigma}(\mathbf{u}_{c}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}\,dS&=0,\end{split}$ (20)
where the second term on the LHS represents the convective transport, and the last term is a source term that accounts for the mass transfer across the interface, which is null when $\dot{m}=0$. Converting the surface integral to a volume integral, we can write it in differential form,
@@
Eq.˜21 is then used to compute the volume fraction, see Section˜2.5.
@@ 2.4 Concentration transport equation
This model can be used to compute the concentration field of the soluble hydrogen in a two-phase system with mass transfer by applying the two-scalar method of fleckenstein2015. In the present study, we focus on pure incompressible gas bubbles, and we assume that no electrolyte species exists in the gas phase (Phase (i.e., the electrolyte is not volatile). The system contains two species: hydrogen (denoted by subscript 1) and electrolyte liquid (denoted by subscript 2). Overall mass transfer is entirely governed by the transport of the hydrogen species. It is worth pointing out that we do not need to solve the mass balance in the dispersed phase since no mixture exists inside the bubbles (pure hydrogen gas). The mass balance of hydrogen in the continuous phase domain $\Omega_{c}$ reads
$\partial_{t}\rho^{1}+\nabla\cdot(\rho^{1}\mathbf{u^{1}})=0,$ (22)
where $\rho^{1}$ is the partial density of hydrogen and $\mathbf{u^{1}}$ denotes the hydrogen velocity. It is coupled with the jump condition for the conservation of mass
$\|\rho^{1}(\mathbf{u^{1}}-\mathbf{u_{\Sigma}})\cdot\mathbf{n_{\Sigma}}\|=\|\dot{m}^{1}\|=0.$ (23)
The average phase density and velocity are derived from the respective species terms
@@
where the superscript $2$ denotes the electrolyte species. Then the transport equation of hydrogen for incompressible flow can be written as
$\partial_{t}\rho^{1}+\mathbf{u}\cdot\nabla\rho^{1}+\nabla\cdot J^{1}=0,$ (26)
@@
The mass-transfer rate of hydrogen can be derived from Eqs˜23 and 14
$\begin{split}\dot{m}^{1}&=\rho^{1}(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}+\rho^{1}(\mathbf{u}^{1}-\mathbf{u})\cdot\mathbf{n}_{\Sigma}\\ &=\frac{\rho^{1}}{\rho_{c}}\cdot\dot{m}+J^{1}\cdot\mathbf{n}_{\Sigma},\end{split}$ (28)
@@
where the $D^{1}$ is the hydrogen diffusion coefficient. Combining Eqs (29) and (28), the mass-transfer rate of hydrogen reads
$\dot{m}=-\frac{M^{1}D^{1}}{1-\frac{\rho^{1}}{\rho_{c}}}\frac{\partial c^{1}}{\partial n_{\Sigma}},$ (30)
where the molar concentration has been introduced, i.e. $c^{1}=\rho^{1}/M^{1}$ with $M^{1}$ the molar mass. It provides the mass-transfer rate evaluated from the continuous side of the interface, which corresponds to $n_{\Sigma}=1$. Since only hydrogen can be transferred between phases, we will omit the species indicator in what follows. We will only refer to the concentration of soluble hydrogen in the liquid phase domain $\Omega_{c}$.
To compute the concentration gradient, we need the hydrogen concentration at the liquid side of the interface. For a gas-liquid system at equilibrium, we can employ Henry’s law to compute the concentration on the liquid side of the interface
$(c_{c})_{\Sigma}=\frac{(c_{d})_{\Sigma}}{P},$ (31)
where $P$ is the partition coefficient, which can be deduced from $P=1/({\mathcal{R}T_{0}H_{s}^{cp}})$, and is taken as a constant for the present work. So, the hydrogen concentration at the liquid side of the interface $(c_{c})_{\Sigma}$ is immediately computed. $(c_{d})_{\Sigma}=\rho_{d}/M$ is a constant, as the density should be constant everywhere inside the bubble ($\Omega_{d}$).
For numerical integration, we rewrite the mass balance Eq.˜26 for the hydrogen diffusion process. As we discussed, the mass balance will be done in the continuous region $\Omega_{c}$ and for a non-reactive flow ($R=0$). Since the hydrogen dissolved in $\Omega_{c}$ is the only species that we need for mass transportation
$\begin{split}\int_{V}\partial_{t}c\,dV&+\oint_{\partial V}(c\mathbf{u}-D\nabla c)\cdot\mathbf{n}\,dS\\ &+\int_{\Sigma}c(\mathbf{u}-\mathbf{u}_{\Sigma})\cdot\mathbf{n}_{\Sigma}\,dS\\ &=0,\end{split}$ (32)
The phase indicator will be omitted in the following, i.e., $c=c_{c}$. Combine the jump condition Eq.˜23, and the final differential form could be deduced,
@@ 2.5 Numerical methodology
The phase change model is derived from the work of Gennari2022. The governing equation shown in Section˜2.3 is solved using the free open science platform Basilisk (HTTP://basilisk.fr/), which provides finite-volume partial differential equation (P.D.E.) solvers on adaptive cartesian grids. Using quadtree/octree adaptive mesh refinement (AMR) in regions with large gradients makes the approach particularly suitable for multiscale processes such as interfacial flows. In interfacial flows, a fine mesh is typically required around the gas-liquid interface but not the entire domain. The shape of the domain is always a square $L_{0}\times L_{0}$ in the axisymmetric configuration (a cube $L_{0}\times L_{0}\times L_{0}$ in three dimensions). The grid is organized following a hierarchical quadtree/octree structure, where each cell can be further divided into four child cells (eight in three dimensions), and a level is assigned to each cell according to its position in the tree structure. The root cell is at level 0, and its size $\Delta$ is the same as that of the whole numerical domain ($\Delta=L_{0}$). A generic cell at level $l$ has size $\Delta(l)=\frac{L_{0}}{2^{l}}$. The grid structure in Basilisk allows neighboring cells to vary by up to one level, meaning each cell edge/face can communicate with no more than two finer edges/faces. The VOF method is one of the most widely used numerical approaches for the modeling of two-phase immiscible fluids. The starting point for the derivation of the VOF approach is the one-fluid formulation presented in Section˜2.3, which is vital to the phase change problem. The volume fraction of the continuous phase is defined as
$f_{c}$ $=\frac{1}{V}\int_{V}H\,dV,$ (34)
@@
The volume fraction of the dispersed phase $f_{d}$ is implicitly described by the relationship $f_{c}+f_{d}=1$. So, only one transport equation of H needs to be solved. By applying the incompressibility constraint, the integrated form of Eq.˜21 is
$\begin{split}\frac{\partial}{\partial t}\int_{V}H\,dV&+\frac{1}{V}\int_{V}\nabla\cdot(H\mathbf{u})\,dV\\ &+\frac{1}{V}\int_{V}\frac{\dot{m}}{\rho_{c}}\delta_{\Sigma}\,dV=0.\end{split}$ (36)
The integration of the transport Eq.˜36 is performed in two steps, namely the reconstruction step and the propagation step. First, the interface is approximated with a line/plane in each interfacial cell. Second, the fluxes of volume fraction across the cell boundaries are computed, and Eq.˜36 is integrated in time. The geometric reconstruction of the interface is based on the piecewise linear interface construction (PLIC) method, where the interface is approximated as a line (plane) in an axisymmetric (three-dimensional) configuration.